ADNAN MUHAMMED

Grade 12,

If (a 2 ,a-2) be a point interior to the region of the parabola y 2 =2x bounded by the chord joining the points (2,2) and (8,-4) then find the set of all possible real values of a.

If (a2,a-2) be a point interior to the region of the parabola y2=2x bounded by the chord joining the points (2,2) and (8,-4) then find the set of all possible real values of a.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
2and y=1/a20 check with respect to origin point lies opposite side of origin with repect to lineput here x=a-4ax2equation of line passing through these two points will be 3y=x+4now here points must lies in the area between parabola and line that meansy